### v1

# On the optimality of MCMC integration for functions with unbounded second moment

Abstract
We study the Markov chain Monte Carlo (MCMC) estimator for numerical integration for functions that do not need to be square integrable w.r.t. the invariant distribution. For chains with a spectral gap we show that the absolute mean error for $L^{p}$ functions, with $p\in(1,2)$, decreases like $n^{1/p-1}$, which is known to be the optimal rate. This improves currently known results where an additional parameter $\delta>0$ appears and the convergence is of order $n^{(1+\delta)/p-1}$.
Keywords: Markov chain Monte Carlo, spectral gap, absolute mean error
Classification: 65C05; 60J22; 65C20
1 Introduction
Let $(\mathcal{X},\mathcal{F}_{\mathcal{X}},\pi)$ be a probability space and $f\colon\mathcal{X}\to\mathbb{R}$ measurable as well as $\pi$-integrable. For a random variable $X\sim\pi$ we are interested in approximating the expectation
$\mathbb{E}[f(X)]=\int_{\mathcal{X}}f(x)\pi({\rm d}x)=\pi(f).$
A common approach is to use an MCMC method. Requiring the density of $\pi$ only in non-normalised form, many MCMC algorithms provide powerful tools in scientific and statistical applications. The main idea behind these approaches is to construct a Markov chain $(X_{n})_{n\in\mathbb{N}}$, having $\pi$ as the invariant distribution, and to estimate $\pi(f)$ via
$S_{n}f=\frac{1}{n}\sum_{j=1}^{n}f(X_{j}).$
Under fairly mild conditions one has $S_{n}f\to\pi(f)$ almost surely as $n\to\infty$, cf. [AG11] or [MT12, Chapter 17]. This ensures (strong) consistency of the MCMC estimator, yet, it is clearly of interest to have non-asymptotic error bounds. For instance, given some $p\in[1,\infty)$, one may consider the $p$-mean error
$\mathbb{E}\left[\left|S_{n}f-\pi(f)\right|^{p}\right],$ (1)
however, also other criteria are feasible, see e.g. [KR19].
Setting $p=2$ in (1) we speak of the mean squared error, and in different settings explicit bounds are known, e.g. under a Wasserstein contraction assumption [JO10], spectral gap conditions [Rud12], or if one has geometric/polynomial ergodicity [ŁMN13]. For the mean squared error to make sense we require a finite second moment of $f$, i.e. $\left\|f\right\|_{L^{2}(\pi)}^{2}=\pi(f^{2})<\infty$. On the other hand, the absolute mean error, given by $\mathbb{E}|S_{n}f-\pi(f)|$, is well defined and finite as long as $f$ is $\pi$-integrable. Bounds for the absolute mean error for functions with $\left\|f\right\|_{L^{p}(\pi)}^{p}=\pi(|f|^{p})<\infty$, where $p<2$, are still rare. In [RS15] it is shown that under a spectral gap condition for any $p\in(1,2)$ one has
$\sup_{\left\|f\right\|_{L^{p}(\pi)}\leq 1}\mathbb{E}[|S_{n}f-\pi(f)|]\leq\frac{C}{n^{1-(1+\delta)/p}},$ (2)
with constants $\delta>0$ and $C\in(0,\infty)$.
Proposition 1 of Section 2.2.9 (with $k=0$) in [Nov88], see also [Hei94, Section 5], shows that in general we have the following lower bound
$\sup_{\left\|f\right\|_{L^{p}(\pi)}\leq 1}\mathbb{E}[|S_{n}f-\pi(f)|]\geq\frac{c}{n^{1-1/p}},$ (3)
where $c\in(0,\infty)$ is a constant independent of $n$.
Even though $\delta>0$ in (2) may be chosen arbitrarily small, it is natural to ask whether it can be removed completely, such that one would have the same rate as in (3). Under the (strong) assumption of uniform ergodicity and reversibility we know that this is the case, cf. [RS15, Theorem 1]. To the best of our knowledge this is the only situation where optimal rates are known. The goal of this note is to extend that result to the spectral gap setting, see Corollary 2.4, where we show
$\sup_{\left\|f\right\|_{L^{p}(\pi)}\leq 1}\mathbb{E}[|S_{n}f-\pi(f)|]\leq\frac{\widetilde{C}}{n^{1-1/p}},$
with an explicit expression for the constant $\widetilde{C}$.
Let us sketch the proof: The main idea is to employ the Riesz-Thorin interpolation theorem, a technique which goes back at least to [Hei94] in Monte Carlo theory and was also used to derive (2) in [RS15]. In contrast to [RS15] we first show a bound for $L^{2}$ functions, see Proposition 3.2, which leads to the optimal rate after interpolation. It is worth mentioning that Proposition 3.2 extends results of [Rud12], see Remark 3.3.
The rest of this note is organised as follows: In Section 2 we state and discuss our assumptions as well as main results. The proofs, together with some intermediate results, can be found in Section 3.
2 Error bounds for MCMC integration
This section contains our main results, see Theorem 2.3 and Corollary 2.4, together with all required notation and assumptions. Let us start with specifying the general setting.
We assume that the state space $(\mathcal{X},\mathcal{F}_{\mathcal{X}})$ is Polish with $\mathcal{F}_{\mathcal{X}}$ being countably generated. Let $K$ be a Markov kernel and $\nu$ a probability measure, called the initial distribution, both defined on $(\mathcal{X},\mathcal{F}_{\mathcal{X}})$. Then, the Markov chain corresponding to $K$ and $\nu$, say $(X_{n})_{n\in\mathbb{N}}$, is defined on a probability space $(\Omega,\mathcal{F},\mathbb{P}_{\nu})$. (In particular, such a probability space exists.) We assume that $\pi$ is the unique invariant distribution of $K$ and that $(X_{n})_{n\in\mathbb{N}}$ is $\psi$-irreducible. For definitions and further details we refer to [MT12, DMPS18].
Given $p\in[1,\infty)$, we define $L^{p}(\pi)$ as the set of all (equivalence classes [^1] Two functions are considered as equivalent if they are equal $\pi$ almost everywhere. of) functions $f\colon\mathcal{X}\to\mathbb{R}$ such that $\left\|f\right\|_{L^{p}(\pi)}<\infty$. Similar, $L^{\infty}(\pi)$ denotes the set of (equivalence classes of) functions with finite $\left\|f\right\|_{L^{\infty}(\pi)}=\operatornamewithlimits{ess\,sup}_{x\in\mathcal{X}}|f(x)|$. Recall that $L^{2}(\pi)$ (equipped with $\langle f,g\rangle_{L^{2}(\pi)}=\int_{\mathcal{X}}f(x)g(x)\pi({\rm d}x)$) is a Hilbert space and so is the following (closed) subspace
$L^{2}_{0}(\pi):=\left\{f\in L^{2}(\pi)\colon\pi(f)=0\right\}.$
Let $p\in[1,\infty]$, then the Markov kernel $K$ induces a linear operator $K\colon L^{p}(\pi)\to L^{p}(\pi)$ via $f\mapsto Kf(\cdot)=\int_{\mathcal{X}}f(x^{\prime})K(\cdot,{\rm d}x^{\prime})$. Indeed, (the operator) $K$ is well defined and one has $\left\|K\right\|_{L^{p}(\pi)\to L^{p}(\pi)}=1$, we refer to [Rud12, Section 3.1] for further details.
[…]
Assumption 2.1.
Assume that $\operatorname{\text{Id}}-K$, considered as operator from $L^{2}_{0}(\pi)$ to $L^{2}_{0}(\pi)$, has a linear and bounded inverse with
$\left\|(\operatorname{\text{Id}}-K)^{-1}\right\|_{L^{2}_{0}(\pi)\to L^{2}_{0}(\pi)}\leq s<\infty.$
Remark 2.2.
The invertibility of $\operatorname{\text{Id}}-K$, restricted to a suitable subspace of $L^{2}(\pi)$, was also studied in [Mat99] for uniformly ergodic chains and [Mat04] for $V$-uniformly ergodic chains. In particular, the existence of $(\operatorname{\text{Id}}-K)^{-1}$ on an appropriate subspace was used there to characterise the convergence behaviour of the mean squared error.
We note that Assumption 2.1 is closely related to a spectral gap, there are, however, different definitions for a spectral gap:
  • Some authors, see e.g. [DMPS18, ALPW23], say $K$ admits a(n) (absolute $L^{2}$) spectral gap if $\sup_{\lambda\in\mathcal{S}_{0}}|\lambda|<1$, where $\mathcal{S}_{0}$ is the spectrum of $K\colon L^{2}_{0}(\pi)\to L^{2}_{0}(\pi)$. This is equivalent to the existence of some $m\in\mathbb{N}$ such that $\left\|K^{m}\right\|_{L^{2}_{0}(\pi)\to L^{2}_{0}(\pi)}<1$, cf. [DMPS18, Proposition 22.2.4].
  • On the other hand, a different definition, for instance used in [Rud12, HSV14, RS15, NRS21], is to say that $K$ admits a(n) (absolute $L^{2}$) spectral gap if $\left\|K\right\|_{L^{2}_{0}(\pi)\to L^{2}_{0}(\pi)}<1$.
If $K$ is reversible (which implies that the corresponding Markov operator is self-adjoint on $L^{2}(\pi)$), then both definitions are equivalent.
Let us emphasize that either of the above definitions of a spectral gap implies that Assumption 2.1 is true, also in the non-reversible case. Spectral gap results were established for a number of MCMC methods, see for instance [HSV14, NRS21, ALPW23], see also [Rud12, Section 3.4] and [RR97, Theorem 2.1]. Moreover, we note that under Assumption 2.1 we cover the setting of Theorems 1 and 2 of [RS15].
Theorem 2.3.
Let Assumption 2.1 be true, $p\in(1,2)$ and assume that $\nu$ is absolutely continuous w.r.t. $\pi$ with Radon-Nikodym derivative $\frac{{\rm d}\nu}{{\rm d}\pi}\in L^{\infty}(\pi)$. Then, for any $f\in L^{p}(\pi)$ and any $n\in\mathbb{N}$ we have that
$\mathbb{E}_{\nu}\left[\left|S_{n}f-\pi(f)\right|^{p}\right]\leq{C_{1}}\frac{\left\|f\right\|_{L^{p}(\pi)}^{p}}{n^{{p-1}}},$
where $C_{1}=\left(1+\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}\right)^{2-p}\left(12s^{2}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}\right)^{p-1}$.
The former results together with Hoelder’s inequality implies the following estimate for the absolute mean error where the rate of convergence is optimal.
Corollary 2.4.
Let Assumption 2.1 be true, $p\in(1,2)$ and assume that $\nu$ is absolutely continuous w.r.t. $\pi$ with Radon-Nikodym derivative $\frac{{\rm d}\nu}{{\rm d}\pi}\in L^{\infty}(\pi)$. Then, for any $f\in L^{p}(\pi)$ and any $n\in\mathbb{N}$ we have that
$\mathbb{E}_{\nu}\left[\left|S_{n}f-\pi(f)\right|\right]\leq{C_{2}}\frac{\left\|f\right\|_{L^{p}(\pi)}}{n^{{1-\frac{1}{p}}}},$
where $C_{2}=\left(1+\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}\right)^{2/p-1}\left(12s^{2}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}\right)^{1-1/p}$.
Remark 2.5.
We note that Theorem 2.3 and Corollary 2.4 are still true (with the same bounds) if we replace $S_{n}f$ by $S_{n,n_{0}}f=\frac{1}{n}\sum_{j=1}^{n}f(X_{j+n_{0}})$, which corresponds to using a burn in of length $n_{0}\in\mathbb{N}$.
Remark 2.6.
In Theorem 2.3 and Corollary 2.4 the quantity $\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}$ appears. In some sense this penalises our choice of the initial distribution $\nu$ which is allowed to differ from the target $\pi$. In the setting where a fixed computational budget is available it may be worth spending some effort to find a “good” initial distribution $\nu$. However, discussing optimal choices of $\nu$ w.r.t. different theoretical and/or practical aspects is beyond the scope of this note.
3 Proofs
In this section we prove our main results. We start with an auxiliary lemma which relates the Monte Carlo sum $S_{n}f$ to a martingale. To do so we use the invertibility of $\operatorname{\text{Id}}-K$ to compute solutions of Poisson’s equation. See also [MT12, Chapter 17.4.1] for further details about that topic.
For any $n\in\mathbb{N}$ we set $\mathcal{F}_{n}=\sigma(X_{j}\colon 1\leq j\leq n+1)$.
Lemma 3.1.
Let Assumption 2.1 hold true and let $h\in L^{2}_{0}(\pi)$. Then, setting $\hat{h}=(\operatorname{\text{Id}}-K)^{-1}h$, we have that $\mathbb{P}_{\nu}$-almost surely
$S_{n}h=\frac{1}{n}M_{n}+\frac{1}{n}\left(\hat{h}(X_{1})-\hat{h}(X_{n+1})\right),$
with $M_{n}=\sum_{j=1}^{n}\left(\hat{h}(X_{j+1})-K\hat{h}(X_{j})\right)$ being an $(\mathcal{F}_{n})_{n\in\mathbb{N}}$ martingale. For its differences $\Delta_{j}=\hat{h}(X_{j+1})-K\hat{h}(X_{j})$ we have
$\forall j\in\mathbb{N}\colon\quad\mathbb{E}_{\nu}\left[|\Delta_{j}|^{2}\right]\leq 2s^{2}\left\|h\right\|_{L^{2}(\pi)}^{2}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}.$
Proof.
Set $T=\operatorname{\text{Id}}-K$. Note that by assumption $\hat{h}\in L^{2}_{0}(\pi)$ and
$h=TT^{-1}h=T\hat{h}=\hat{h}-K\hat{h}.$
Thus $\mathbb{P}_{\nu}$-almost surely
$S_{n}h$ $=\frac{1}{n}\sum_{j=1}^{n}\hat{h}(X_{j})-K\hat{h}(X_{j})$
$=\frac{1}{n}\sum_{j=1}^{n}\left(\hat{h}(X_{j+1})-K\hat{h}(X_{j})\right)+\frac{1}{n}\left(\hat{h}(X_{1})-\hat{h}(X_{n+1})\right)$
$=\frac{1}{n}M_{n}+\frac{1}{n}\left(\hat{h}(X_{1})-\hat{h}(X_{n+1})\right).$
Since $(X_{n})_{n\in\mathbb{N}}$ is a Markov chain with kernel $K$ for any $n\geq 2$ holds
$\mathbb{E}_{\nu}[M_{n}|\mathcal{F}_{n-1}]=M_{n-1}+\mathbb{E}_{\nu}\left[\hat{h}(X_{n+1})-K\hat{h}(X_{n})\;\big{|}\;\mathcal{F}_{n-1}\right]=M_{n-1}.$
It remains to show the bound for the differences $\Delta_{j}$, which also implies that $\mathbb{E}_{\nu}|M_{n}|<\infty$ for any $n\in\mathbb{N}$ and hence that $(M_{n})_{n\in\mathbb{N}}$ is a martingale. Thereto observe that $\|\hat{h}\|_{L^{2}(\pi)}\leq s\left\|h\right\|_{L^{2}(\pi)}$ and that $\left\|K\right\|_{L^{1}(\pi)\to L^{1}(\pi)}=1$. Hence, Hoelder’s inequality shows that for any $j\in\mathbb{N}$ holds
$\mathbb{E}_{\nu}\left[|\hat{h}(X_{j+1})|^{2}\right]=\int_{\mathcal{X}}K^{j}|\hat{h}|^{2}(x)\frac{{\rm d}\nu}{{\rm d}\pi}(x)\pi({\rm d}x)\leq\left\|K^{j}|\hat{h}|^{2}\right\|_{L^{1}(\pi)}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}$
$\leq\|\hat{h}\|_{L^{2}(\pi)}^{2}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}\leq s^{2}\|h\|_{L^{2}(\pi)}^{2}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}.$
Similarly, $\mathbb{E}_{\nu}[(K\hat{h}(X_{j}))^{2}]\leq\mathbb{E}_{\nu}[K\hat{h}^{2}(X_{j})]\leq s^{2}\|h\|_{L^{2}(\pi)}^{2}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}$. The bound for $\mathbb{E}_{\nu}|\Delta_{j}|^{2}$ now follows by using the inequality $(\alpha+\beta)^{2}\leq 2(\alpha^{2}+\beta^{2})$, being valid for any $\alpha,\beta\in\mathbb{R}$. ∎
Based on the above martingale decomposition of Lemma 3.1 we now show a bound for the mean squared error.
Proposition 3.2.
Let Assumption 2.1 be true and assume that $\nu$ is absolutely continuous w.r.t. $\pi$ with Radon-Nikodym derivative $\frac{{\rm d}\nu}{{\rm d}\pi}\in L^{\infty}(\pi)$. Then, for any $f\in L^{2}(\pi)$ we have that
$\mathbb{E}_{\nu}\left[\left|S_{n}f-\pi(f)\right|^{2}\right]\leq 6s^{2}\left(1+\frac{1}{n}\right)\frac{\left\|f\right\|_{L^{2}(\pi)}^{2}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}}{n}.$
Proof.
Let $h=f-\pi(f)\in L^{2}_{0}(\pi)$. Lemma 3.1 and the inequality $(\alpha+\beta+\gamma)^{2}\leq 3(\alpha^{2}+\beta^{2}+\gamma^{2})$, which is true for any real numbers $\alpha,\beta,\gamma$, yield
$\mathbb{E}_{\nu}\left[|S_{n}f-\pi(f)|^{2}\right]=\mathbb{E}_{\nu}\left[(S_{n}h)^{2}\right]\leq\frac{3}{n^{2}}\mathbb{E}_{\nu}\left[M_{n}^{2}\right]+\frac{3}{n^{2}}\mathbb{E}_{\nu}[\hat{h}(X_{1})^{2}+\hat{h}(X_{n+1})^{2}],$
with $M_{n},\hat{h}$ as specified in Lemma 3.1. By the martingale property (and again using the notation/results of Lemma 3.1) it follows that
$\mathbb{E}_{\nu}\left[M_{n}^{2}\right]=\mathbb{E}_{\nu}\left[\sum_{j=1}^{n}\Delta_{j}^{2}\right]\leq\sum_{j=1}^{n}\mathbb{E}_{\nu}\left[\Delta_{j}^{2}\right]\leq 2ns^{2}\|h\|_{L^{2}(\pi)}^{2}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}.$
Furthermore, the same arguments as in the proof of Lemma 3.1 reveal
$\mathbb{E}_{\nu}[\hat{h}(X_{1})^{2}+\hat{h}(X_{n+1})^{2}]\leq 2s^{2}\|h\|_{L^{2}(\pi)}^{2}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}.$
Combining these estimates with $\left\|h\right\|_{L^{2}(\pi)}=\left\|f\right\|_{L^{2}(\pi)}^{2}-\pi(f)^{2}\leq\left\|f\right\|_{L^{2}(\pi)}^{2}$ yield the desired result. ∎
Remark 3.3.
Note that Proposition 3.2 can be seen as extension of two results in [Rud12]:
  • It extends Theorem 3.34 to the case where $K$ is not required to be reversible [^2] Note that $L^{1}$-exponential convergence implies $L^{2}$-exponential convergence, e.g. by an application of the Riesz-Thorin interpolation theorem, see [Gra14, Theorem 1.3.4]. Thus Assumption 2.1 holds true in this case..
  • It extends Theorem 3.41 to the $L^{2}$ setting.
Finally, we turn to the proof of our main result.
Proof of Theorem 2.3.
We write $L^{q}(\mathbb{P}_{\nu})$ for the set of random variables $X$ with finite $\left\|X\right\|_{L^{q}(\mathbb{P}_{\nu})}^{q}=\mathbb{E}_{\nu}|X|^{q}$. For $n\in\mathbb{N}$ consider the linear mapping $T_{n}f=S_{n}f-\pi(f)$ and note that
$\mathbb{E}_{\nu}|T_{n}f|\leq\frac{1}{n}\sum_{j=1}^{n}\mathbb{E}_{\nu}|f(X_{j})|+\pi(|f|)=\frac{1}{n}\sum_{j=1}^{n}\int_{\mathcal{X}}K^{j-1}|f|(x)\frac{{\rm d}\nu}{{\rm d}\pi}(x)\pi({\rm d}x)+\pi(|f|).$
By Hoelder’s inequality and the fact that $\left\|K\right\|_{L^{1}(\pi)\to L^{1}(\pi)}=1$ it follows that $\left\|T_{n}\right\|_{L^{1}(\pi)\to L^{1}(\mathbb{P}_{\nu})}\leq 1+\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}$.
Additionally, by Proposition 3.2 holds
$\left\|T_{n}\right\|_{L^{2}(\pi)\to L^{2}(\mathbb{P}_{\nu})}\leq\frac{1}{\sqrt{n}}\left(12s^{2}\left\|\frac{{\rm d}\nu}{{\rm d}\pi}\right\|_{L^{\infty}(\pi)}\right)^{1/2}.$
An application of the Riesz-Thorin interpolation theorem, see [Gra14, Theorem 1.3.4] in the setting $p_{0}=q_{0}=1$; $p_{1}=q_{1}=2$ and $\theta=2-2/p$, yields that
$\left\|T_{n}\right\|_{L^{p}(\pi)\to L^{p}(\mathbb{P}_{\nu})}\leq\left\|T_{n}\right\|_{L^{1}(\pi)\to L^{1}(\mathbb{P}_{\nu})}^{1-\theta}\left\|T_{n}\right\|_{L^{2}(\pi)\to L^{2}(\mathbb{P}_{\nu})}^{\theta},$
which, together with the above estimates, implies the result. ∎
Acknowledgements
The author would like to thank Mareike Hasenpflug and Daniel Rudolf for valuable discussions which helped to significantly improve the presentation of the paper. JH is supported by the DFG within project 432680300 – SFB 1456 subproject B02.